Reports 1-1 of 1 Clear search Modify search
AdV-DAQ (Data Acquisition and Global Control)
masserot - 22:36 Wednesday 03 March 2021 (50973) Print this report
TSC HWS NI,WI Temperature control setup - digital part

The temperature probes are acquired at 100Hz by the ADC7674  SN43 labelled as  "TCS_ADC_Moni2"  located into the TSC room using the following channels, the bias and the slope remains to be tuned

  • ADC7674_CHANNEL TCS_HWS_NI_TE1                  2    100        Butterworth_8_20dB            0.                +1                "V"
  • ADC7674_CHANNEL TCS_HWS_NI_TE2                  3    100        Butterworth_8_20dB            0.                +1                "V"
  • ADC7674_CHANNEL TCS_HWS_WI_TE1                  4    100        Butterworth_8_20dB            0.                +1                "V"
  • ADC7674_CHANNEL TCS_HWS_WI_TE2                  5    100        Butterworth_8_20dB            0.                +1                "V"

These channels are sent to the SUSP_rtpc(rtpc4) and made available for 2 dedicated Acl servers: TCS_HWS_NI and TCS_HSW_WI  running at 100Hz

These servers have been setup using a configuration already tested in the Virgo RD VPM domain with few minors modifiactions:

  • an ACL_RELAY_CH called ENBL has been added to  enable the loop and to reset the filters when the loop is opened
  • the filter computation method has been set to use the long double computation to reduce the possible digital noise
  • a include file with the filters definition  has been created to ease the filters modification
  • the correction channel CORR_U has been made available as 64 bits double for the server TCS_CEB mananing the DAC1955 SN40

The TCS_CEB server, running at 10KHz,

  • read the TCS_HWS_{NI,WI}_CORR_U (line25-26)
  • extend the channels at the DAC1955 running frequency(100KHz)
  • implement  a clipping (line 72-73) at 0 and 1V  after the channel extension  to be tuned if needed


Comments to this report:
masserot - 11:41 Tuesday 23 March 2021 (51182) Print this report

Today the setups for NE and WE HWS Temperature control have been implemented

using the hardware informations in the logbook 51010

Search Help